OpenStudy (anonymous):

How does log base 2 (x+1)(x+5)=4 turn into: 2^4= (x+1)(x-5) ?

OpenStudy (dumbcow):

thats how the log is defined \[\large a^x = y \rightarrow x = \log_{a} y\]

OpenStudy (anonymous):

log(base 2 ) (x +1)(x + 5) = 4 [log (x+1)(x+5)]/log 2 = 4 [log (x+1)(x+5)] = 4log2 [log (x+1)(x+5)] = log(2^4) taking antilog on both sides, (x+1)(x+5) = 2^4

OpenStudy (anonymous):

@mitul why would you divide by log 2? can't you only do that with the quotient property? and i don't think we have learned antilogs yet...

OpenStudy (anonymous):

dividing by log 2 is due to a property called change of base.. whenever we have log(base a)b, we can write as log(base exponential)b/log(base exponential)a or simply ln(b)/ln(a). ln means natural log.

OpenStudy (anonymous):

And antilog is just the opposite property of log just as positive nullifies negative!!

OpenStudy (anonymous):

Okay thank you....do you know of a loop method because we learned that in school but i don't quite understand it...?

OpenStudy (anonymous):

i don't know the name but i just might be able to identify the method if u can show me wat is it??

OpenStudy (anonymous):

this is what we learned for a simpler one and we are supposed to apply the same principal i think...  thats what we learned in class